# Active Light Modulation to Counter Manipulation of Speech Visual Content

### v1

@@ A.1. Adaptive Optical Embedding
$I$ : updated required intensity for SLM cells.
@@ 15            $c^{j}_{SLM}\leftarrow$ Equation1($c^{j}_{p_{off}},I^{j}$)
Algorithm 1 Adaptation performed after completion of a window to set upcoming SLM bitmap cell colors and intensities.
The below equation, employed on Line 1 configures bitmap colors to match the patch of projection surface they are incident upon.
@@
subj. to $\alpha*R^{j}_{p_{off}}+\alpha*G^{j}_{p_{off}}+\alpha*B^{j}_{p_{off}}=I_{j},$
@@
The user-specified values of $\phi_{max}$ and $\beta_{max}$ configure the tradeoff between data error rate, in the form of BER and perceptibility. $\delta$ tunes the system’s adaptation speed, with a higher value enabling faster adaptation but risking ”overshooting” in terms of perceptibility. Furthermore, we note that if BER is above $\beta_{max}$, all $I^{j}$ are incremented by $\delta$. The rationale for the global increase in Line 1 is that, while cells’ perceptibilities are independent, robustness depends on all cells since data is distributed across them. We also note $I^{j}$ may be incremented even when CIEDE2000$(c_{p_{on}}^{j},c_{p_{off}}^{j})$ exceeds $\Phi_{max}$, consciously prioritizing robustness. This is a conscious design decision to prioritize robustness, as we believe the impact of an unrecoverable signature to be worse than that of embedding visibility.
A.2. Optical Modulation Configuration

### v2

@@ 5.2. Spatio-Temporal Light Modulation
II : updated required intensity for SLM cells.
@@ 15    $c^{j}_{SLM}\leftarrow$ Equation1($c^{j}_{p_{off}},I^{j}$)
Algorithm 1 Adaptation performed after completion of a window to set upcoming SLM bitmap cell colors and intensities. $\phi_{max}$ and $\beta_{max}$ configure the trade-off between robustness and imperceptibility. $\delta$ controls the adaptation speed. A higher $\delta$ value corresponds to a more aggressive response to BER spikes.
5.3. Adaptive Embedding
Thus, we propose an intensity-guided adaptive embedding method (Algorithm 1), which continually adapts the cell intensities $I^{j}$ required for sufficient SNR and then optimizes cell colors $c_{SLM}^{j}$ under this constraint. Prior to VeriLight’s deployment, we perform a short, one-time calibration enabling the core unit to map SLM bitmap pixels to its camera’s pixels. Upon completion of a window, VeriLight runs Algorithm 1 in parallel with ongoing modulations. It assesses robustness and perceptibility based on the completed windows’ video, and accordingly increments or decrements $I^{j}$. To quantify robustness, the core unit runs data extraction on its past window’s video (§6) and computes the bit error rate (BER). The perceptibility of each cell $j$ is measured as the perceived difference between patch $j$’s color with and without SLM light, approximated via the CIEDE2000 color difference formula (Sharma et al., 2005). Then, $c_{SLM}^{j}$ is chosen to reduce perceptibility while satisfying the intensity requirement $I_{j}$:
@@ 5.3. Adaptive Embedding
subj. to $\alpha*R^{j}_{p_{off}}+\alpha*G^{j}_{p_{off}}+\alpha*B^{j}_{p_{off}}=I^{j},$